A solution contains 0.159 M C2H5NH3 and 0.177 M ethylamine, C2H5NH2.

The pH of this solution is

.
Answer #1

Kb Kb 10-17(43) = 9.37 931ナ1090-59 9:32 P 14-932 H= H-68